I do not understand how to solve equations. eample
-2x+8=104+6x
People have tried to explain to me and when I think I have it figured out, it's wrong. I need a simple way of solving these types of equations.. "HELP"

Answer by Alan3354(69443)

-2x+8=104+6x
--------------------
ok, try something different
get everything on one side.
-2x+8=104+6x
Subtract what's on the right from both sides.
-2x+8 -(104+6x) = 0
Multiply thru by the -1 (the minus sign)
-2x+8 -104-6x = 0
add the x's, and the numbers:
-8x - 96 = 0
-----------------
Now get an x on one side by itself.
-8x - 96 = 0
get rid of the 96 by adding 96 to both sides
-8x = 96
Divide both sides by -8
x = 96/-8
x = -12
Done